OpenStudy (anonymous):

1. The Grand Prix Formula 1 race this weekend is in Singapore. The race consists of 61 laps around a 5.067km long track. At 225 miles per hour, how long will it take to go around once?

OpenStudy (nincompoop):

how many miles are there in a kilometer or vice-versa?

OpenStudy (tmak11):

1: 0.621371 according to google

OpenStudy (tmak11):

the track is 3.148486857 miles aroud

OpenStudy (tmak11):

i know im wrong but i keep ketting 0.01399327492 hours

OpenStudy (tmak11):

well that got complicated fast

OpenStudy (nincompoop):

it doesn't it follows proper conversion

OpenStudy (tmak11):

1lap is roughly 0.84 min or 50.375789712 seconds

OpenStudy (nincompoop):

\[\sf \frac{5.067 km}{61 laps} \rightarrow \approx \frac{8.31 \times 10^{-2} km}{1 lap} \times \frac{1mi}{6.21 \times 10^{-1}km} \times \frac{1 hr}{225 mi} \times \frac{60min}{1hr}\approx \frac{xmin}{lap}\]

OpenStudy (nincompoop):

units cancel and convert properly

OpenStudy (nincompoop):

kilometers cancel out, miles cancel out, then the hour is converted into minutes so basically you are left with minutes per lap unit

OpenStudy (nincompoop):

the question is how long will it take to go around once that is hour/lap

OpenStudy (nincompoop):

or min/lap or sec/lap depends on what you want, but it should be time/lap